OpenStudy (anonymous):

Simplify 6 – 4x – 2 + x. A. 4 + x B.1 C. 4 - 3x D. 3 - 4x

Mehek (mehek14):

6 - 4 = ? -4x + x = ?

OpenStudy (anonymous):

6 - 4 = 2 but i'm still not getting the second part :/

OpenStudy (ddos_dragon):

The answer is C BECAUSE -

Mehek (mehek14):

-4x + x = -4 + 1 = ?

OpenStudy (ddos_dragon):

6 – 4x – 2 + x. Take the +1x and add it to the -4X Get -3x

OpenStudy (ddos_dragon):

Take the 6 -2 get 4

OpenStudy (ddos_dragon):

4 - 3X

OpenStudy (ddos_dragon):

Tis' C.

OpenStudy (ddos_dragon):

Remember to medal, as well as close the question.

OpenStudy (anonymous):

Thank you! That was correct @Ddos_Dragon